# DustNET: enabling machine learning and AI models of dusty plasmas

### 23 Mar 2026 v2

@@ III.2 Dust models
In reality, dust charging remains an open problem for practical reasons and and there is no ‘universal formula’ that can be used to calculate grain charge accurately for all situations. Besides unique material and structural properties, such as dust electric conductivity (dielectric or conducting), surface conditions such as absorption, material phases (liquid or solid), more complicated dust shapes other than an ideal spherical or cylindrical geometry, plasma parameters and properties, such as electron density, Debye length, ion physics, kinetic effects, dust charging is best dealt with numerical methods. First-principle methods, such as Quantum MD simulations Dunjko and Briegel (2018), involve solution of Schrödinger’s equation but are not often used in dusty plasmas. Tthe time characteristics of dust charging, such as the relaxation time to reach the steady grain charge and the charge fluctuations of grains of different sizes Maiorov et al. (2002), also needs to be analyzed numerically on a case-by-case basis. As an example, Fig. 10, by employing OML theory with a line-of-sight (LOS) modification, the charging currents to the surface of an aggregate grain can be calculated Matthews et al. (2007); Matthews and Hyde (2007); Matthews et al. (2012). The code is also adapted to model the charging of aggregates or individual dust grains which are in close proximity by removing any charging currents which are blocked by a nearby grain.
@@
Another force classification scheme is based on the force scaling with dust dimensions. divided into body-type forces ($\propto r_{d}^{3}\sim V_{d}$, an example is gravity), surface-type forces ($\propto r_{d}^{2}\sim S_{d}$, examples include neutral drag force, ion drag force, and laser pressure force) or size-type forces ($\propto r_{d}\sim l_{d}$, an example is electric force) Fortov et al. (2005); Piel (2017b). Gravitational force in a strongly coupled dusty plasma, new attractive forces such as dipole-dipole interactions and wakefield may also Shukla (2001). The dust particles interact through Coulomb forces mediated by the surrounding plasma Melzer and Goree (2008); Morfill and Ivlev (2009), and as a result the effective forces between particles can be non-reciprocal and break energy conservation Melzer et al. (2019); Nikolaev and Timofeev (2021); Kolotinskii et al. (2021); Vaulina et al. (2015); Ivlev et al. (2015).
@@ III.4 Model reduction and acceleration
Recent efforts have focused on data-driven reduced models, when the governing equation of the evolution may be unknown, particularly those through the use of neural networks Kim et al. (2023). For instance, the relatively simple artificial neural network (ANN), in today’s standard, was trained to reproduce the spectral energy distributions predicted by a hybrid code comprised of the GALFORM semi-analytical model of galaxy formation, which predicts the full star formation and galaxy merger histories, and the GRASIL spectro-photometric code, which carries out a self-consistent calculation of the spectral energy distributions, including absorption and emission of radiation by dust Almeida et al. (2010). The ANN had a feed-forward neural network architecture with 12 input neurons, one hidden layer with 60 neurons, or two hidden layers with 30 neurons each, and an output layer with one or more neurons. More recently, techniques like autoencoders, neural ODEs, and physics-informed neural networks (PINNs) are being applied to extract low-dimensional dynamics from simulation or experimental data Padula et al. (2024).
@@ III.5 Machine learning & AI models
The model assumed that the horizontal ($xy$ plane) acceleration of each particle was determined by the boxed equation in Fig. 13, where $\vec{\rho}_{i}=(x_{i},y_{i})$, $\vec{\rho}_{ij}=(x_{i}-x_{j},y_{i}-y_{j})=\rho_{ij}\hat{\rho}_{ij}$, and $f_{ij}=F_{ij}/m_{i}$, ($F_{ij}$ is the magnitude of the interaction force). The reduced environmental force was $\vec{f}^{\text{env}}_{i}=\vec{F}^{\text{env}}_{i}/m_{i}$, where $\vec{F}^{\text{env}}_{i}$ is the horizontal environmental force on particle $i$, and the damping coefficient of particle $i$ was $\gamma_{i}$. In the experiments, the ion wakes beneath each particle resulted in non-reciprocal effective forces ($F_{ij}\neq F_{ji}$, Fig. 13A). Furthermore, the particles were not identical, and required particle-level identifiers. Ideally, this would be the mass of each particle, which is unknown. Yet heavier particles sit lower in the plasma sheath, and thus the mean $z$-position, averaged over an entire time series: $s_{i}=\left<z_{i}\right>_{t}$, was used as a good identifier for each particle. For each of the forces, a separate, fully-connected NN was used to map the input positions and identifiers to their respective forces, and the 3 NNs were trained in parallel (Fig. 13D). To learn each force accurately, the inputs to each term contained minimal overlap (i.e., environmental forces depend on $x_{i}$ and $y_{i}$, but interaction forces only depend on the relative distance $\rho_{ij}$). The loss function compared the measured acceleration of each particle to the model acceleration given by the 3 NNs, and was summed over all time points.
@@
Yet major challenges remain. First, one needs to formalize the notion of “sufficient but not too restrictive” physical constraints. Then it would be possible to automate the design and application of NN models to plasma experiments, so that various engineering choices do not depend on the intuition of a particular researcher. The second challenge is learning field-level descriptions of dusty plasma. Dusty plasmas often have more particles than can be tracked accurately, and particles can be 100 nm or less (too small to individually resolve). For this challenge, one would need to combine methods that can identify parsimonious continuum equations Brunton et al. (2016); Champion et al. (2019); Alves and Fiuza (2022); Reinbold et al. (2021); Gurevich et al. (2019) with those that extract low-dimensional predictive embeddings (aka, relevant variables) from data LeCun (2022); Abdelaleem et al. (2023), and then to endow both with the useful physical intuition.
@@ IV.2 Dusty plasmas in fusion devices
A Dust Ablation Trail Analysis (DATA) code, a deep-learning enhanced package, is being developed to automatically detect and analyze dust ablation trails from experiment images captured in tokamak discharges Liang et al. (2023b). Due to the limited viewing access to tokamaks such as the Experimental Advanced Superconducting Tokamak (EAST), it is not practical to obtain a sufficiently large experimental datasets of dust ablation trails with different dust types, orientations, intensity, and sizes. Leveraging the similarities between plasma jets and fusion dust ablation trails in the shape and size, laboratory plasma jet images were used to train and validate the DATA code package Liang et al. (2023b). A plasma jet inside a quartz tube was produced from a steady argon gas flow at a rate varying from 3 and $6$ standard liters per minute. Thousands of plasma jet images have been obtained using two cameras by varying relative orientation of the jets to the cameras, relative distances, and the plasma jet size in a controlled setting.
@@
Next, the Zhang-Suen thinning algorithm was used in DATA to quantitatively characterize the classified plasma jets zhang1984fast. The principle of the Zhang-Suen thinning algorithm is to reduce the area of the object image without changing its topology, so that the skeleton information of the object is achieved. The skeleton information can be used for the further investigation of plasma jets, such as the three-dimensional (3D) reconstruction.
@@ IV.3 Space microgravity dusty plasmas
Machine-learning techniques have been used to identify and quantify structures from (2D) video images. Examples include the identification of “dust strings” (dust particles that become aligned due to ion flows in the plasma) Dietz et al. (2021). In PK-4 rapid switching of the electric field polarity within the plasma leads to the formation of a string-like dust fluid, a so-called electrorheological plasma Ivlev et al. (2008); Dietz et al. (2021). Other example applications of AI and ML include the identification of hexatic phases from the 2D arrangement of particles Du et al. (2024) or the identification of the solid-liquid transition line from supervised and unsupervised learning Li et al. (2024a). In addition, for simulated data, it was possible to locally assign 3D crystal structures in a dust system to fcc, hcp, and bcc crystal structures or disordered particles Dietz et al. (2017). This approach requires the input of the 3D particle positions.
@@
ML and AI have been envisioned for experimental data processing and controls in ISS. The Zyflex chamber Knapek et al. (2021), being the heart of the COMPACT experiment on the ISS and currently undergoing parabolic flight testing led by the University of Greifswald, is equipped with a center and a ring electrode both on the upper and lower side. The electrodes are driven by a four-channel rf-generator that allows to independently apply rf signals with given amplitude and relative phase. In a pulsed rf-mode, it is possible to confine extended and quite homogeneous dust clouds without central void Knapek et al. (2024) which is a prerequisite for most of the experiments planned with COMPACT, see Fig. 18.
@@ IV.6 Phase transition: Melting of two-dimensional dusty plasmas
Deep learning, particularly convolutional neural network (ConvNet), has been widely applied in the image analysis. As the experimental results in dusty plasma research are often recorded as sequence of images, it is natural to apply ConvNet in the data analysis. Lack of experimental data poses challenges in the application of deep learning. To address this issue, Langevin dynamics simulations have been performed. The equation of motion including the damping from the neutral gas and Brownian motion of microparticles is given by
$m\ddot{\bm{r}}_{i}+m\nu\dot{\bm{r}}_{i}=-\sum_{j\neq i}\bigtriangledown\phi_{ij}+\bm{L}_{i},$ (5)
where $\bm{r}_{i}$ is the position of particle $i$, $m$ is the particle mass, and $\nu$ is the damping rate, which results from the neutral gas. The Brownian motion was included in the equation and the corresponding Langevin force $\bm{L}_{i}$ at a certain kinetic temperature $T$ was defined by $\left\langle\bm{L}_{i}\right\rangle=0$ and $\left\langle\bm{L}_{i}(t)\bm{L}_{j}(t+\tau)\right\rangle=2\nu mk_{b}T\delta_{i,j}\delta(\tau)\bm{I}$, where $\delta_{ij}$ is Kronecker delta, $\delta(\tau)$ is the delta function, and $\bm{I}$ is the unit matrix. In the simulation, it is commonly assumed that particles interact with each other via Yukawa potential Ivlev et al. (2011).
To study the melting process of plasma crystal, a heating process is performed in the simulation, as demonstrated in Fig. 24(a). Initially, the system is equilibrated for a certain amount of time at a low temperature and then gradually heated at a small rate, until the plasma crystal is fully melted. The local structure around particle $i$ could be quantified by the hexatic order parameter $\Psi_{6,i}=\frac{1}{6}\sum_{k=1}^{6}{e^{j6\theta_{k}}}$, where six nearest neighboring particles were considered and $\theta_{k}$ is the angle between $\bm{r}_{k}-\bm{r}_{i}$ and the $x$ axis. If particle $i$ is located in the center of a perfect hexagon cell, its $\Psi_{6}$ is unity, as shown in the left inset in Fig. 24(a).
@@
For the purpose of model training and later application in the experiment analysis, sequences of images are prepared based on the simulation results. The particle positions at each time step in the simulation are converted to a gray-scale bitmap image, resembling the image obtained from the video recording in the experiments. Particularly, the intensity distribution of each bright spot in the synthesized image, corresponding to a particle, is assumed to follow the Gaussian distribution. The size is generally much larger than the real size of the particle, due to the diffraction by the particle as well as camera properties such as diffraction by the camera aperture and imperfect lens focusing. However, in the experiments, due to the presence of overexposure and noise in the recording, the brightness distribution of each particle deviates from the Gaussian distribution to some extent. To reduce such discrepancy, a threshold is applied to the such images to remove noise, binarize the gray-scale bitmaps, and convolve the resulting binary images with the Gaussian filter. These steps ensure that the algorithm trained with the simulation results can be directly applied to the experiment analysis Huang et al. (2022b); Li et al. (2024a); Du et al. (2024).
ConvNet is employed to find the melting line. ReLU activation function are used for each convolutional layer and the softmax activation function is used for the outpuut layer. Two machine learning strategies can be applied to investigate the melting transition and identify the melting line. In the supervised learning method, the training samples in the form of synthesized images at very high and low coupling are labeled, corresponding to the crystalline and liquid phases. Samples with a thermodynamic state near the critical value are not provided to the training, to avoid ambiguity Carrasquilla and Melko (2017); Li et al. (2021). Once the training is completed, the model is applied to classify the complete data set covering the full range of the coupling parameter. As the bitmap images are fed to the ConvNet model, it returns a probability $P(\Gamma,N)$ where the images are classified as plasma crystal. Here $N$ is the approximate number of particles included in the bitmap image sample and is selected as $N=80$ for the realization. The crossing of $P(\Gamma,N)$ and $1-P(\Gamma,N)$ provides the coupling strength of the melting transition $\Gamma_{m}$.
@@
Figure 26: Identification of the hexatic phase with convolutional neural network. (a) Probability of dusty plasma being classified as crystalline, hexaitc, and liquid and (b) concentrations of dislocation pairs, dislocations, and disclinations as function of reduced kinetic temperature $T/T_{m}$ in the experiment and simulation. Reproduced from Du et al. (2024).
@@
The ConvNet is easy to deploy using an open source platform for machine learning and efficient in application. For example, for a training sample containing $2000$ labeled gray-scale bitmap images with a size of $100\times 100$ pixel2, it takes $5$ minutes to train a model with batches of $100$ images for $20$ epochs on a desktop machine with Intel(R) Core i7-10900 CPU and takes seconds to classify an image from experiments. This is much more efficient than the scheme using traditional particle tracking and order parameter analysis. Such methods can also be extended to classify the liquid and crystalline structure in three-dimensional dusty plasmas in the laboratory and in the payload on board the Space Station Thomas et al. (2008); Pustylnik et al. (2016); Knapek et al. (2021).

### 09 Apr 2026 v3

@@ III.2 Dust models
In reality, dust charging remains an open problem for practical reasons and there is no ‘universal formula’ that can be used to calculate grain charge accurately for all situations. Besides unique material and structural properties, such as dust electric conductivity (dielectric or conducting), surface conditions such as absorption Bronold and Fehske (2016); Mengel et al. (2024), material phases (liquid or solid), more complicated dust shapes other than an ideal spherical or cylindrical geometry Matyash and Schneider (2006); Ikkurthi et al. (2008), plasma parameters and properties, such as electron density, Debye length, ion physics, kinetic effects, dust charging is best dealt with numerical methods. First-principle methods, such as Quantum MD simulations Dunjko and Briegel (2018), involve solution of Schrödinger’s equation but are not often used in dusty plasmas. The time characteristics of dust charging, such as the relaxation time to reach the steady grain charge and the charge fluctuations of grains of different sizes Maiorov et al. (2002), also needs to be analyzed numerically on a case-by-case basis. As an example, Fig. 10, by employing OML theory with a line-of-sight (LOS) modification, the charging currents to the surface of an aggregate grain can be calculated Matthews et al. (2007); Matthews and Hyde (2007); Matthews et al. (2012). The code is also adapted to model the charging of aggregates or individual dust grains which are in close proximity by removing any charging currents which are blocked by a nearby grain.
@@
Another force classification scheme is based on the force scaling with dust dimensions, as summarized in Fig. 11: body-type forces ($\propto r_{d}^{3}\sim V_{d}$, an example is gravity), surface-type forces ($\propto r_{d}^{2}\sim S_{d}$, examples include neutral drag force, ion drag force, and laser pressure force) or size-type forces ($\propto r_{d}\sim l_{d}$, an example is electric force) Fortov et al. (2005); Piel (2017b). New classes of effective forces, with a size-scaling of $r_{d}^{\gamma}$ and the exponent $\gamma$ being a non-integer, may also arise. Examples may include dipole-dipole interactions and wakefield effects Shukla (2001), Coulomb forces modified by the surrounding plasma Melzer and Goree (2008); Morfill and Ivlev (2009). As a result, the effective forces between particles can be non-reciprocal and break energy conservation Melzer et al. (2019); Nikolaev and Timofeev (2021); Kolotinskii et al. (2021); Vaulina et al. (2015); Ivlev et al. (2015).
@@ III.4 Model reduction and acceleration
Recent efforts have focused on data-driven reduced models, when the governing equation of the evolution may be unknown, particularly those through the use of neural networks Kim et al. (2023). For instance, the relatively simple artificial neural network (ANN), in today’s standard, was trained to reproduce the spectral energy distributions predicted by a hybrid code comprised of the GALFORM semi-analytical model of galaxy formation, which predicts the full star formation and galaxy merger histories, and the GRASIL spectro-photometric code, which carries out a self-consistent calculation of the spectral energy distributions, including absorption and emission of radiation by dust Almeida et al. (2010). The ANN had a feed-forward neural network architecture with 12 input neurons, one hidden layer with 60 neurons, or two hidden layers with 30 neurons each, and an output layer with one or more neurons. More recently, techniques like autoencoders, neural ODEs Lu et al. (2021), and PINNs are being applied to extract low-dimensional dynamics from simulation or experimental data Padula et al. (2024).
@@ III.5 Machine learning & AI models
The model assumed that the horizontal (xyxy plane) acceleration of each particle was determined by the boxed equation in Fig. 13, where $\vec{\rho}_{i}=(x_{i},y_{i})$, $\vec{\rho}_{ij}=(x_{i}-x_{j},y_{i}-y_{j})=\rho_{ij}\hat{\rho}_{ij}$, and $f_{ij}=F_{ij}/m_{i}$, ($F_{ij}$ is the magnitude of the interaction force). The reduced environmental force was $\vec{f}^{\text{env}}_{i}=\vec{F}^{\text{env}}_{i}/m_{i}$, where $\vec{F}^{\text{env}}_{i}$ is the horizontal environmental force on particle $i$, and the damping coefficient of particle $i$ was $\gamma_{i}$. In the experiments, the ion wakes beneath each particle resulted in non-reciprocal effective forces ($F_{ij}\neq F_{ji}$, Fig. 13A). Furthermore, the particles were not identical, and required particle-level identifiers. Ideally, this would be the mass of each particle, which is unknown. Yet heavier particles sit lower in the plasma sheath, and thus the mean $z$-position, averaged over an entire time series: $s_{i}=\left<z_{i}\right>_{t}$, was used as a good identifier for each particle. For each of the forces, a separate, fully-connected NN was used to map the input positions and identifiers to their respective forces, and the 3 NNs were trained in parallel (Fig. 13D). To learn each force accurately, the inputs to each term contained minimal overlap (i.e., environmental forces depend on $x_{i}$ and $y_{i}$, but interaction forces only depend on the relative distance $\rho_{ij}$). The loss function compared the measured acceleration of each particle to the model acceleration given by the 3 NNs, and was summed over all time points.
@@
Yet major challenges remain. First, one needs to formalize the notion of “sufficient but not too restrictive” physical constraints. Then it would be possible to automate the design and application of NN models to plasma experiments, so that various engineering choices do not depend on the intuition of a particular researcher. The second challenge is learning field-level descriptions of dusty plasma. Dusty plasmas often have more particles than can be tracked accurately, and particles can be 100 nm or less (too small to individually resolve). For this challenge, one would need to combine methods that can identify parsimonious continuum equations Brunton et al. (2016); Champion et al. (2019); Alves and Fiuza (2022); Reinbold et al. (2021); Gurevich et al. (2019) with those that extract low-dimensional predictive embeddings (aka, relevant variables) from data LeCun (2022); Abdelaleem et al. (2023), and then to endow both with the useful physical intuition. A more detailed discussion of specific neural network architectures and their scaling with data size is presented in Sec. V.
@@ IV.2 Dusty plasmas in fusion devices
A Dust Ablation Trail Analysis (DATA) code, a deep-learning enhanced package, is being developed to automatically detect and analyze dust ablation trails from experiment images captured in tokamak discharges Liang et al. (2023b). Due to the limited viewing access to tokamaks such as the Experimental Advanced Superconducting Tokamak (EAST), it is not practical to obtain a sufficiently large experimental datasets of dust ablation trails with different dust types, orientations, intensity, and sizes. Leveraging the similarities between plasma jets and fusion dust ablation trails in the shape and size, laboratory plasma jet images were used to train and validate the DATA code package Liang et al. (2023b). A plasma jet inside a quartz tube was produced from a steady argon gas flow at a rate varying from 3 and 66 standard liters per minute. Thousands of plasma jet images have been obtained using two cameras by varying relative orientation of the jets to the cameras, relative distances, and the plasma jet size in a controlled setting.
@@
Next, the Zhang-Suen thinning algorithm was used in DATA to quantitatively characterize the classified plasma jets Zhang and Suen (1984). The principle of the Zhang-Suen thinning algorithm is to reduce the area of the object image without changing its topology, so that the skeleton information of the object is achieved. The skeleton information can be used for the further investigation of plasma jets, such as the three-dimensional (3D) reconstruction.
@@ IV.3 Space microgravity dusty plasmas
Machine-learning techniques have been used to identify and quantify structures from (2D) video images. Examples include the identification of “dust strings” (dust particles that become aligned due to ion flows in the plasma) Dietz et al. (2021). In PK-4 rapid switching of the electric field polarity within the plasma leads to the formation of a string-like dust fluid, a so-called electrorheological plasma Ivlev et al. (2008); Dietz et al. (2021). Other example applications of AI and ML include the identification of hexatic phases from the 2D arrangement of particles Du et al. (2024) or the identification of the solid-liquid transition line from supervised and unsupervised learning Li et al. (2024a), see Sec. IV.6 for further discussions. In addition, for simulated data, it was possible to locally assign 3D crystal structures in a dust system to fcc, hcp, and bcc crystal structures or disordered particles Dietz et al. (2017). This approach requires the input of the 3D particle positions.
@@
ML and AI have been envisioned for experimental data processing and controls aboard the ISS. The Zyflex chamber Knapek et al. (2021), being the heart of the COMPACT experiment on the ISS and currently undergoing parabolic flight testing led by the University of Greifswald, is equipped with a center and a ring electrode both on the upper and lower side. The electrodes are driven by a four-channel rf-generator that allows to independently apply rf signals with given amplitude and relative phase. In a pulsed rf-mode, it is possible to confine extended and quite homogeneous dust clouds without central void Knapek et al. (2024) which is a prerequisite for most of the experiments planned with COMPACT, see Fig. 18.
@@ IV.6 Phase transition: Melting of two-dimensional dusty plasmas
Deep learning, particularly convolutional neural network (ConvNet), has been widely applied in the image analysis. As the experimental results in dusty plasma research are often recorded as sequence of images, it is natural to apply ConvNet in the data analysis. Lack of experimental data poses challenges in the application of deep learning. To address this issue, MD (aka, Langevin dynamics) simulations have been performed. To study the melting process of plasma crystal, a heating process is performed in the simulation, as demonstrated in Fig. 24(a). Initially, the system is equilibrated for a certain amount of time at a low temperature and then gradually heated at a small rate, until the plasma crystal is fully melted. The local structure around particle $i$ could be quantified by the hexatic order parameter $\Psi_{6,i}=\frac{1}{6}\sum_{k=1}^{6}{e^{j6\theta_{k}}}$, where six nearest neighboring particles were considered and $\theta_{k}$ is the angle between $\bm{r}_{k}-\bm{r}_{i}$ and the $x$ axis. If particle $i$ is located in the center of a perfect hexagon cell, its $\Psi_{6}$ is unity, as shown in the left inset in Fig. 24(a).
@@
For the purpose of model training and later application in the experiment analysis, sequences of images are prepared based on the simulation results. The particle positions at each time step in the simulation are converted to a gray-scale bitmap image, resembling the image obtained from the video recording in the experiments. Particularly, the intensity distribution of each bright spot in the synthesized image, corresponding to a particle, is assumed to follow the Gaussian distribution. The size is generally much larger than the real size of the particle, due to the diffraction by the particle as well as camera properties such as diffraction by the camera aperture and imperfect lens focusing. However, in the experiments, due to the presence of overexposure and noise in the recording, the brightness distribution of each particle deviates from the Gaussian distribution to some extent. To reduce such discrepancy, a threshold is applied to the images to remove noise, binarize the gray-scale bitmaps, and convolve the resulting binary images with the Gaussian filter. These steps ensure that the algorithm trained with the simulation results can be directly applied to the experiment analysis Huang et al. (2022b); Li et al. (2024a); Du et al. (2024).
ConvNet is employed to find the melting line, Fig. 25. ReLU activation function is used for each convolutional layer and the softmax activation function is used for the outpuut layer. Two machine learning strategies can be applied to investigate the melting transition and identify the melting line. In the supervised learning method, the training samples in the form of synthesized images at very high and low coupling are labeled, corresponding to the crystalline and liquid phases. Samples with a thermodynamic state near the critical value are not provided to the training, to avoid ambiguity Carrasquilla and Melko (2017); Li et al. (2021). Once the training is completed, the model is applied to classify the complete data set covering the full range of the coupling parameter. As the bitmap images are fed to the ConvNet model, it returns a probability $P(\Gamma,N)$ where the images are classified as plasma crystal. Here NN is the approximate number of particles included in the bitmap image sample and is selected as $N=80$ for the realization. The crossing of $P(\Gamma,N)$ and $1-P(\Gamma,N)$ provides the coupling strength of the melting transition $\Gamma_{m}$.
@@
Figure 26: Identification of the hexatic phase with convolutional neural network. (a) Probability of dusty plasma being classified as crystalline, hexatic, and liquid and (b) concentrations of dislocation pairs, dislocations, and disclinations as function of reduced kinetic temperature $T/T_{m}$ in the experiment and simulation. Reproduced from Du et al. (2024).
@@
The ConvNet is easy to deploy using an open source platform for machine learning and efficient in application. For example, for a training sample containing $2000$ labeled gray-scale bitmap images with a size of $100\times 100$ pixel2, it takes 55 minutes to train a model with batches of $100$ images for 2020 epochs on a desktop machine with Intel(R) Core i7-10900 CPU and takes seconds to classify an image from experiments. This is much more efficient than the scheme using traditional particle tracking and order parameter analysis. Such methods can also be extended to classify the liquid and crystalline structure in three-dimensional dusty plasmas in the laboratory and in the payload on board the Space Station Thomas et al. (2008); Pustylnik et al. (2016); Knapek et al. (2021).